OpenStudy (anonymous):

Rationalize the denominator. Assume all expressions under radicals represent positive numbers. 3sqrt5y^4 over 3sqrt10x^4

OpenStudy (anonymous):

\[3\sqrt 5y^4 \over 3\sqrt10x^4\]

OpenStudy (anonymous):

There we go. That's what I was looking for. :) What does it mean to rationalize the denominator?

OpenStudy (anonymous):

no clue

OpenStudy (anonymous):

I would assume to find the common den.

myininaya (myininaya):

if you want \[\sqrt{10x^4}\] you can write sqrt{10x^4} in equation editor just so you know

OpenStudy (anonymous):

Well, the square root of a non-perfect square is known as an irrational number, so right now the denominator is an irrational number. We want to "rationalize" it by getting rid of that square root.

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

so we just have 10x^4?

OpenStudy (anonymous):

And is it just the square root of 10 or the expression 10x^4

OpenStudy (anonymous):

k

myininaya (myininaya):

is it \[\sqrt{10}x^4 or \sqrt{10x^4}\]

OpenStudy (anonymous):

Thanks myin.

OpenStudy (anonymous):

the second one. I messed up

OpenStudy (anonymous):

Okay, so what can I multiply that by to get rid of the square root? Any ideas or thoughts?

myininaya (myininaya):

and i assume its the second way for the numerator well except with different numbers

OpenStudy (anonymous):

2?

myininaya (myininaya):

anyways we want to rationalize the denominator? so are there any perfect squares under the square root?

OpenStudy (anonymous):

I'll give you a hint. The square root of x is the same as x^(1/2) If I take x^(1/2) * x^(1/2) what do I get?

OpenStudy (anonymous):

Remember what we talked about with exponents on our last problem! :)

OpenStudy (anonymous):

0

OpenStudy (anonymous):

We are multiplying this time; not dividing.

OpenStudy (anonymous):

1

OpenStudy (anonymous):

Yes, 1. So if I took \[\sqrt{x} \times \sqrt{?}\]

OpenStudy (anonymous):

Oops. I meant to write the square root of x both times.

OpenStudy (anonymous):

x^2?

OpenStudy (anonymous):

If I multiply the square root of x by the square root of x I got x, didn't I ?

OpenStudy (anonymous):

so it's just x?

OpenStudy (anonymous):

Remember x^(1/2) * x^(1/2) = x^(1/2 + 1/2) = x^1 = x

OpenStudy (anonymous):

Yes, it's just x. So if I have the square root of 10x^4, what can I multiply that by to get rid of the square root?

OpenStudy (anonymous):

-4 idk i'm getting confused again lol

OpenStudy (anonymous):

4

OpenStudy (anonymous):

Ok. \[\sqrt{10x^4} = (10x^4)^{1/2}\]

OpenStudy (anonymous):

So how could I make sure the exponent equals 1?

OpenStudy (anonymous):

no idea = / my luck, it's probably really simple to figure out

OpenStudy (anonymous):

Remember x^(1/2) * x^(1/2) = x So (10x^4)^(1/2) * (????)^(1/2) = 10x^4

OpenStudy (anonymous):

1

OpenStudy (anonymous):

Basically, if you have the square root of a value and you multiply it by the square root of that same value, you ALWAYS get the value of what is inside the radical.

OpenStudy (anonymous):

\[\sqrt{10x^4} \times \sqrt{10x^4} = 10x^4\]

OpenStudy (anonymous):

It works for anything.

OpenStudy (anonymous):

ooooh. ok.

OpenStudy (anonymous):

It's because you are adding the exponents which are 1/2 and 1/2 + 1/2 = 1

OpenStudy (anonymous):

So back to the original question, you would multiply both the numerator and the denominator by sqrt(10x^4)

OpenStudy (anonymous):

The first 2 steps are actually easier than this. Can we cancel out the 3s? Can we do anything with the square root of the variables?

OpenStudy (anonymous):

cancel them out as well?

OpenStudy (anonymous):

We can cancel out the 3s. However, if we look at x^4 that's the same as x * x * x * x. How many groups of 2 Xs are there?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

\[\sqrt{x^4} = x^2\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Exactly! So we can rewrite this as the following: \[y^2\sqrt{5} / x^2\sqrt{10}\]

OpenStudy (anonymous):

So now all we need to do is multiply the numerator and denominator by what to get rid of that sqrt(10) ?

OpenStudy (anonymous):

x^2sqrt10?

OpenStudy (anonymous):

Close. Just sqrt(10)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The x^2 is fine on its own

OpenStudy (anonymous):

So if you multiply the top and bottom by that, what are you left with?

OpenStudy (anonymous):

5y and x^4?

OpenStudy (anonymous):

y^2 * sqrt(5) * sqrt(10) / x^2 * sqrt(10) * sqrt(10) y^2 * sqrt(50) / x^2 * 10

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

\[y^2\sqrt{5}\sqrt{10} = y^2\sqrt{50}\]

OpenStudy (anonymous):

\[x^2\sqrt{10}\sqrt{10} = 10x^2\]

OpenStudy (anonymous):

Can you simplify sqrt(50) any further?

OpenStudy (anonymous):

sqrt10?

OpenStudy (anonymous):

50 = 2 x 5 x 5

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So I can pull out a 5 and am left with \[5\sqrt{2}\]

OpenStudy (anonymous):

2*5

OpenStudy (anonymous):

The 5 then will simplify with the 10 in the denominator leaving a 2 in the denominator.

OpenStudy (anonymous):

oooh

OpenStudy (anonymous):

So.. We are left with \[y^2\sqrt{2} / 2x^2\]

OpenStudy (anonymous):

No more irrational stuff in the denominator!

OpenStudy (anonymous):

awesome!

OpenStudy (anonymous):

Yeah, if you're into that sort of thing.